### v1

# Revealing Spin and Spatial Symmetry Decoupling: New Insights into Magnetic Systems with Dzyaloshinskii-Moriya Interaction

Abstract
It is widely accepted that spin-orbit coupling (SOC) generally locks spin and spatial degrees of freedom, as a result, the spin, despite being an axial vector, is fixed and cannot rotate independently, and the magnetic system should be described by magnetic space groups (MSGs). While as a new type of group, spin space groups (SSGs) have been introduced to approximately describe the symmetry of magnetic systems with negligible SOC, and received significant attention recently. In this work, we prove that in two cases of coplanar spin configurations, there are spin-only operations that strictly hold even with considerable Dzyaloshinskii-Moriya interaction (DMI), and the symmetry of their spin models could be described by the spin-coplanar SSG. In addition, we also find that for spin-collinear cases, regardless the strength of DMI, the magnon systems within the framework of linear spin wave theory (LSWT) also preserve the decoupled spin and spatial rotations, but the symmetry does not belong to the conventional definitions of collinear spin groups. Finally, we discuss the potential realization of these novel symmetries for magnetic candidate materials in rod, layer, and three-dimensional (3D) space groups. Our work extends the applicability of SSGs to magnetic materials with heavy elements, and provides new avenues for exploring novel physical phenomena in magnon topology and transport.
Symmetry, as formulated by group theory, is fundamental to the laws of nature. In condensed-matter physics, symmetry has long been central, as it dictates the physical properties, phase transitions, and behavior of various materialsDresselhaus et al. (2007). It governs the conservation laws Baym and Kadanoff (1961), determines selection rules Moskovits (1982), and plays a particularly crucial role in fields of Landau theory of phase transition Landau and Lifshitz (2013), topological phases Chiu et al. (2016); Fu (2011); Zhang et al. (2019); Vergniory et al. (2019); Tang et al. (2019) and quasiparticle excitations Bradlyn et al. (2016); Yu et al. (2022); Tang and Wan (2021). The symmetries of three-dimensional (3D) solids are comprehensively described by crystallographic group theory, with spatial operations, including rotations, reflections, translations, and their combinations, forming the 32 point groups (PGs) and 230 space groups (SGs) for nonmagnetic systems Bradley and Cracknell (2009). When spin-orbit coupling (SOC) effect is considered, the double SGs are applicable. On the other hand, for magnetic materials, the SOC generally locks spin and spatial degrees of freedom, preventing spin from rotating independently, and the symmetry of magnetic systems should be described by 122 magnetic point groups (MPGs) and 1651 magnetic space groups (MSGs) Bradley and Cracknell (2009).
In 1960s, it is found that the Heisenberg model neglecting the Dzyaloshinskii-Moriya interaction (DMI), which is the leading order of SOC effect Dzyaloshinsky (1958); Moriya (1960), enjoys higher symmetries than MSGs, and the concept of spin space groups (SSGs) is introduced Brinkman and Elliott (1966a, b); Litvin and Opechowski (1974); Litvin (1977). It has attracted significant attention in areas such as Landau theory of phase transitions Litvin et al. (1982); Izyumov and Syromyatnikov (2012), neutron scattering Elliott and Thorpe (1967), and magnon band topology Chisnell et al. (2015); Li et al. (2017); Corticelli et al. (2022, 2023). Very recently, the full classification of collinear SSGs and the associated magnon band topology for collinear magnets Chen et al. (2023) are developed, offering new pathways to exploring exotic phenomena of magnonic systems.
In SSGs, the spin and spatial operations become partially decoupled, giving rise to an expanded set of symmetries expressed as Litvin and Opechowski (1974); Litvin (1977); Liu et al. (2022),
$G_{SS}=G_{NSS}\times G_{SO}$ (1)
where $G_{SO}$ stands for the spin-only group that only contains spin operations, $G_{NSS}$ stands for the nontrivial spin space group that contains no pure spin operations Liu et al. (2022).
In recent years, SSGs have been extended to electronic systems, and drawn growing attention due to their relevance in the study of magnetic materials with negligible SOC Liu et al. (2022); Yang et al. (2024); Guo et al. (2021); González-Hernández et al. (2021); Bai et al. (2022); Šmejkal et al. (2022a); Ahn et al. (2019); Hayami et al. (2020); Yuan et al. (2021); Šmejkal et al. (2020); Yuan et al. (2020); Šmejkal et al. (2022b, c); Zhou et al. (2024); Lee et al. (2024); Gomonay et al. (2024); Liao et al. (2024); Bose et al. (2022); Zhu et al. (2024); Krempaskỳ et al. (2024); Autieri (2024); Xiao et al. (2024); Jiang et al. (2024); Chen et al. (2024). In field of magnetic topological electronic states, the classifications and the irreducible corepresentations of spin groups for coplanar and collinear magnetic structures are derived Liu et al. (2022), and the Z2 topological phases Liu et al. (2022) and exotic nodal-line or -point semimetals Yang et al. (2024); Guo et al. (2021) protected by SSG symmetry are predicted. Meanwhile, SSGs are also crucial for understanding the unconventional spin-momentum locking without SOCGonzález-Hernández et al. (2021); Bai et al. (2022); Šmejkal et al. (2022a); Ahn et al. (2019); Hayami et al. (2020); Yuan et al. (2021); Šmejkal et al. (2020); Yuan et al. (2020), which was later recognized as a characteristic feature of altermagnetismŠmejkal et al. (2022b, c); Zhou et al. (2024); Lee et al. (2024); Gomonay et al. (2024); Liao et al. (2024); Bose et al. (2022); Zhu et al. (2024); Krempaskỳ et al. (2024); Autieri (2024). Very recently, an exhaustive analysis of SSGs symmetries provides a full classification and investigates the enriched features of electronic bands within SSGs framework Xiao et al. (2024); Jiang et al. (2024); Chen et al. (2024), which paves the way for deeper understanding and further exploration of emergent phenomena in magnetic materials.
The symmetry of SSGs is generally an approximation, holding strictly only when SOC is absent. However, as a relativistic effect, SOC always exists and plays a crucial role in various fields of condensed matter physics, such as permanent magnets Campbell (1996); Skomski and Coey (2019), spintronics Bader and Parkin (2010), spin-orbitronics Manchon et al. (2015), and topological physics Hasan and Kane (2010); Qi and Zhang (2011). The DMI, which has been widely regarded as the “SOC” in spin models and magnon systems Brinkman and Elliott (1966a); Onose et al. (2010); Cheng et al. (2016); Zyuzin and Kovalev (2016); Chumak et al. (2015); McClarty (2022), is also always present even in the centrosymmetric compound, since not all bond centers are inversion centers. In topological magnons, DMI generates the Berry curvature and plays a central role in the study of magnon band topology Zhang et al. (2013); Mook et al. (2016); Su et al. (2017); Kim et al. (2016); Wang et al. (2018); Su et al. (2017); Owerre (2017). Meanwhile, in the field of magnon transport, DMI is essential in many phenomena such as the magnon Hall effectOnose et al. (2010); Di et al. (2015); Hirschberger et al. (2015); Matsumoto and Murakami (2011); Doki et al. (2018); Lee et al. (2015); Mook et al. (2019); Laurell and Fiete (2018) and the magnon spin Nernst effectZyuzin and Kovalev (2016); Cheng et al. (2016); Kovalev and Zyuzin (2016); Li et al. (2020); Zyuzin and Kovalev (2018); Chen et al. (2021), where DMI takes on a role analogous to that of SOC in the spin Hall effect.
The DMI and SSG symmetries have both played significant roles in the exploration of exotic phenomena in magnetic materials. However, as mentioned above, they are inherently in conflict. Could there be scenarios where SSG symmetries remain strictly applicable even when huge DMI is taken into account? The coexistence of DMI and SSG symmetries in one magnetic system will undoubtedly provide new avenues for study novel physical phenomena. In this work, we prove that the inclusion of DMI does not necessarily lead to a complete locking of spin and spatial rotations, and identify a novel symmetry to describe the spin model and magnon system with even considerable DMIs.
Group Theory Analysis
First we consider the coplanar spin configurations with the spin moments lying in the $xy$ plane. The symmetry of spin-coplanar SSGs could be written as Eq. (1), where the spin-only group for spin-coplanar SSGs has the expresssion as Litvin and Opechowski (1974); Litvin (1977); Liu et al. (2022),
$G_{SO}=\{E,TU_{z}(\pi)\}$ (2)
where $T$ and $U_{z}(\pi)$ represent time-reversal symmetry and the two-fold spin rotation operation along the z-axis, respectively.
In the spin Hamiltonian, the antisymmetric DMI is expressed as $H_{DMI}=\sum_{i,j}\bm{D_{ij}}\cdot(\bm{S_{i}}\times\bm{S_{j}})$ Dzyaloshinsky (1958); Moriya (1960). In general, the inclusion of DMI locks spin and spatial rotations, and reduces the Hamiltonian symmetry to MSGs. However, we find that under two following cases of magnetic configurations, the spin and spatial rotations are still partially decoupled:
(i) The first case occurs in 2D systems, when the magnetic atoms lie within the $xy$ plane, which possesses horizontal mirror symmetry $\sigma_{h}$. In this case, only the component of DMI perpendicular to the mirror plane, i.e. $D^{z}$, is left Moriya (1960), which makes that the spin Hamiltonian retains the symmetry $\{TU_{z}(\pi)||E\}$. Therefore, the spin-only group $G_{SO}=\{E,TU_{z}(\pi)\}$ in spin-coplanar SSG is strictly preserved. Below we provide a criterion to determine whether the operations in $G_{NSS}$ are preserved:
An operation $R\in G_{NSS}$ in 2D systems could be written as $R=\{A||B\}$, where $A$ and $B$ represents the spin and spatial operations, respectively. The generators for the spin operation $A$ include $C_{nz}$ and $C_{2x}$, where $n=1,2,3,4,6$Liu et al. (2022), while the generators for the spatial operation $B$ include $C_{nz}$, $C_{2x}$, $S_{4}$, $P$, $\sigma_{h}$ and $\sigma_{v}$itvin and Wike (2012), where $S_{4}$ presents the four-fold rotation inversion operation and $P$ presents inversion operation, while $\sigma_{h}$ and $\sigma_{v}$ correspond to mirror operactions along the horizontal and vertical planes, respectively. The DMI with only $D^{z}$ can be expressed as $H_{DMI}=\sum_{i,j}D_{ij}^{z}(S_{i}^{x}S_{j}^{y}-S_{i}^{y}S_{j}^{x})$. By applying the generators to the DMI term, $S_{i}^{x}S_{j}^{y}-S_{i}^{y}S_{j}^{x}$ changes sign under $C_{2x}$ and $\sigma_{v}$, while $C_{nz}$, $S_{4}$, $P$ and $\sigma_{h}$ leaves $S_{i}^{x}S_{j}^{y}-S_{i}^{y}S_{j}^{x}$ invariant. For a given operation $R=\{A||B\}$, when the operations $A$ and $B$ both change or both preserve the sign of $S_{i}^{x}S_{j}^{y}-S_{i}^{y}S_{j}^{x}$, the operation $R$ is preserved even in the presence of DMI. Otherwise, the operation $R$ would be broken.
(ii) The second case arises in 1D systems, when the magnetic atoms are aligned along $z$-axis chain, and the chain has double rotational symmetry $C_{2z}$, which also constrain the direction of the DMI to $z$ directionMoriya (1960). Therefore, the DMI with only $z$-components $D^{z}$ ensures that the symmetry of spin Hamiltonian maintain the form of spin-coplanar SSG, which could be written as Eq. (1) with the spin-only group $G_{SO}=\{E,TU_{z}(\pi)\}$.
In magnon spintronics, the ability to generate a pure transverse spin current without a thermal current is of great interest Cheng et al. (2016); Zyuzin and Kovalev (2016), and we expect that it can be achieved in the above two cases. Generally, DMI is the driving force behind the magnon thermal Hall effect Onose et al. (2010); Di et al. (2015); Hirschberger et al. (2015); Matsumoto and Murakami (2011); Doki et al. (2018). However, in the above two cases, the preserved spin-only operation $\{TU_{z}(\pi)||E\}$ ensures that the magnon Berry curvature is an odd function with respect to the quasimomentum $\bm{k}$, $\Omega_{n}(\bm{k})=-\Omega_{n}(-\bm{k})$, which leads to the absence of magnon thermal Hall response Mook et al. (2019); Laurell and Fiete (2018). While, the pure magnon spin current driven by a temperature gradient can still exist in these two spin-coplanar cases.
Furthermore, for collinear spin configuration within the framework of the linear spin wave theory (LSWT), we also find interesting symmetries here. The spin-only group $G_{SO}$ for the collinear spin configurations could be written as Litvin and Opechowski (1974); Litvin (1977); Liu et al. (2022),
$G_{SO}=\{E,TU_{n}(\pi)\}\ltimes U_{z}(\phi)$ (3)
where the direction of spin moments is defined along the $z$-direction, and $U_{z}(\phi),$ $\phi\in[0,2\pi)$ means full spin rotation along $z$-axis, $\{E,TU_{n}(\pi)\}$ contains the identity operation as well as the combined operation of time reversal $T$ and two-fold spin rotation about $n$-axis perpendicular to the spin $z$-axis. Within the framework of LSWT, the Holstein-Primakoff (HP) transformation results in that only $D_{ij}^{z}$ term contains quadratic contribution of magnons. Through symmetry analysis, we demonstrate that the $D_{ij}^{z}$ term preserves $\{U_{z}(\phi)||E\}$ symmetry but breaking the operation $\{TU_{n}(\pi)||E\}$. Therefore, the magnons in spin-collinear systems has the novel symmetry expressed as $G_{SS}=G_{NSS}\times G_{SO}^{\prime}$, where the spin-only group could be written as
$G_{SO}^{\prime}=U_{z}(\phi)$ (4)
while the nontrivial spin space group $G_{NSS}$ is a subgroup of the one for original SSG without DMI. It should be noted that, the above expression of the spin group $G_{SS}$ has spin-only symmetry beyond MSG, but does not belong to the conventional definitions of collinear or coplanar spin groups, whose spin-only groups are given by Eq. (2) and (3) respectively.
In this scenario, the spin-only group $G_{SO}^{\prime}$ also have significant impact on the physical phenomena in magnon transport. Since $\{U_{z}(\phi)||E\}$ symmetry is preserved, $S^{z}$ remains a good quantum number, thus magnons with opposite chirality do not couple with each another when DMI is taken into account Onose et al. (2010); Cheng et al. (2016); Zyuzin and Kovalev (2016). Meanwhile, the introduction of the DMI breaks the $\{TU_{n}(\pi)||E\}$ symmetry, thereby enabling the non-zero magnon Hall response and magnon spin conductivity, which provides a symmetry-based explanation for why the DMI is crucial in the magnon Hall effect and magnon spin Nernst effect in previous studies Onose et al. (2010); Cheng et al. (2016); Zyuzin and Kovalev (2016).
Table 1: We list all 17 layer groups with mirror symmetry $\sigma_{h}$, if the magnetic atoms of 2D materials occupy the given Wyckoff positions (i.e. $z$=0) and their spins lie within the xy plane, the spin Hamiltonian including DMI maintain the symmetry of spin-coplanar SSG, written as Eq. (1) and Eq. (2).
Layer Group Wyckoff Positions
$p11m$ 1a
$p112/m$ 1a,1b,1c,1d,2i
$pm2m$ 1a,1b,2e
$pb2_{1}m$ 2a
$cm2m$ 2a,4c
$pmmm$ 1a,1b,1c,1d,2e,2f,2g,2h,4q
$pmam$ 2a,2b,2c,4f
$pbam$ 2a,2b,4e
$cmmm$ 2a,2b,4c,4d,4e,8k
$p4/m$ 1a,1b,2c,4g
$p4/mmm$ 1a,1b,2c,4g,4h,4i,8j
$p4/mbm$ 2a,2b,4e,8f
$p-6$ 1a,1b,1c,3g
$p6/m$ 1a,2b,3d,6g
$p-6m2$ 1a,1b,1c,3g,6h
$p-62m$ 1a,2b,3d,6g
$p6/mmm$ 1a,2b,3d,6g,6h,12k
Table 2: The selected 2D candidate materials from C2DB Haastrup et al. (2018); Gjerding et al. (2021), with magnetic atoms occupying given Wyckoff positions in TABLE 1 and exhibiting in-plane spin configurations.
Layer Group Materials
$pm2m$ $\mathrm{NiCl_{2}N_{2}H_{4}C_{6}}$,$\mathrm{MoV_{3}Se_{8}}$,
$\mathrm{WV_{3}Se_{8}}$,$\mathrm{MoV_{3}Te_{8}}$,
$\mathrm{WV_{3}Te_{8}}$,$\mathrm{OVBr_{2}}$,
$\mathrm{OVCl_{2}}$,$\mathrm{OVF_{2}}$,
$\mathrm{OVI_{2}}$,$\mathrm{Mo_{2}V_{2}Se_{8}}$,
$\mathrm{Mo_{2}V_{2}Te_{8}}$,$\mathrm{V_{2}W_{2}Te_{8}}$
$cm2m$ $\mathrm{FeOBr_{2}}$
$pmmm$ $\mathrm{CrOF_{2}}$
$cmmm$ $\mathrm{OsVO_{2}Br_{4}}$
$p4/mmm$ $\mathrm{Co_{2}S_{2}}$,$\mathrm{Ni_{2}S_{2}}$
$p-6m2$ $\mathrm{VS_{2}}$,$\mathrm{VSe_{2}}$,$\mathrm{VTe_{2}}$,
$\mathrm{YBr_{2}}$,$\mathrm{YI_{2}}$,$\mathrm{FeBr_{2}}$,
$\mathrm{FeI_{2}}$,$\mathrm{MnTe_{2}}$,$\mathrm{NbTe_{2}}$,
$\mathrm{ScBr_{2}}$,$\mathrm{ScCl_{2}}$,$\mathrm{ScI_{2}}$,
$\mathrm{ScSe_{2}}$,$\mathrm{CrAs_{2}}$,$\mathrm{CrBi_{2}}$
$p-62m$ $\mathrm{Ti_{2}I_{6}}$
Candidate Materials
Based on the above symmtry analysis, we will extract the criteria to identify the magnetic materials whether the spin and spatial operations are still partially decoupled even with large DMI. The 2D systems are described by wallpaper groups or layer groups Hahn et al. (1983). In the case of wallpaper groups, only the $xy$-direction is defined, thus the $xy$-plane is inherently mirror symmetric. For all the 17 wallpaper groups, when the spins lie in the $xy$-plane, the spin-only group $G_{SO}$ in spin-coplanar SSG is preserved.
Table 3: We provide the symmetry criteria applicable to 3D magnetic systems by mapping 2D layer groups to 3D space groups as discussed in the main text. In these space groups, when magnetic atoms occupy given Wyckoff positions with coplanar magnetic moments, and the interlayer distance is relatively large, the huge DMI does not degrade the spin Hamiltonian symmetry to MSGs, but preserve the expression of SSG symmetry with spin-only group as Eq. (2).
Layer Group Space Group Wyckoff Positions(Space Group)
$p11m$ $Pm$ 1a,1b
$p112/m$ $P2/m$ 1a-1h,2m,2n
$pm2m$ $Pmm2$ 1a-1d,2e-2f
$pb2_{1}m$ $Pmc2_{1}$ 2a,2b
$cm2m$ $Amm2$ 2a,2b,4d,4e
$pmmm$ $Pmmm$ 1a-1h,2i-2p,4y,4z
$pmam$ $Pmma$ 2a-2f,4i,4j
$pbam$ $Pbam$ 2a-2d,4g,4h
$cmmm$ $Cmmm$ 2a-2d,4e-4j,8p,8q
$p4/m$ $P4/m$ 1a-1d,2e,2f,4j,4k
$p4/mmm$ $P4/mmm$ 1a-1d,2e,2f,4j-4o,8p,8q
$p4/mbm$ $P4/mbm$ 2a-2d,4g,4h,8i,8j
$p-6$ $P-6$ 1a-1f,3j,3k
$p6/m$ $P6/m$ 1a,1b,2c,2d,3f,3g,6j,6k
$p-6m2$ $P-6m2$ 1a-1f,3j,3k,6l,6m
$p-62m$ $P-62m$ 1a,1b,2c,2d,3f,3g,6j,6k
$p6/mmm$ $P6/mmm$ 1a,1b,2c,2d,3f,3g,6j-6m,12p,12q
Table 4: We list the 1D rod groups with twofold rotational symmetry $C_{2z}$ along the $z$-axis. With spins lying in the $xy$ plane, the magnetic materials that meet the symmetry criteria outlined in the table will exhibit additional SSG symmetry.
Rod Group Wyckoff Positions Rod Group Wyckoff Positions
$p112$ 1a $p-42m$ 1a,1b,2c
$p112/m$ 1a,1b,2c $p-42c$ 2a,2b,4e
$p222$ 1a,1b, 2g $p4/mmm$ 1a,1b,2c
$pmm2$ 1a $p4/mcc$ 2a,2b,4c
$pcc2$ 2a $p4_{2}/mmc$ 2a,2b,4c
$pmmm$ 1a,1b,2g $p6$ 1a
$pccm$ 2a,2b,4e $p6_{2}$ 3a
$p4$ 1a $p6_{4}$ 3a
$p4_{2}$ 2a $p6/m$ 1a,1b,2c
$p-4$ 1a,1b,2c $p622$ 1a,1b,2c
$p4/m$ 1a,1b,2c $p6_{2}22$ 3a,3b,6c
$p4_{2}/m$ 2a,2b,4c $p6_{4}22$ 3a,3b,6c
$p422$ 1a,1b,2c $p6mm$ 1a
$p4_{2}22$ 2a,2b,4c $p6cc$ 2a
$p4mm$ 1a $p6/mmm$ 1a,1b,2c
$p4_{2}cm$ 2a $p6/mcc$ 2a,2b,4c
$p4cc$ 2a
On the other hand, since most 2D materials typically consist of multiple atomic layers or have finite extension in the $z$-direction, the layer group usually provides a more appropriate description Hahn et al. (1983). There are 17 out of 80 layer groups having the mirror symmetry $\sigma_{h}$, which are summarized in TABLE 1. If a magnetic material belongs to the layer groups in TABLE 1, and its magnetic atoms are also located at the given Wyckoff positions in TABLE 1, i.e. $z=0$, while their spins lie in the $xy$-plane, the presence of DMI does not degrade the spin Hamiltonian symmetry to MSGs, but instead preserves the spin-planar SSGs symmetries.
For any 2D magnetic material, one can easily identify their symmetries using our TABLE 1. To illustrate the result, we apply our theory to a popular database Computational 2D Materials Database (C2DB)Haastrup et al. (2018); Gjerding et al. (2021) and identify 33 materials as summarized in TABLE 2. The spin Hamiltonian for these materials would maintain the SSG symmetry even with large DMI.
In addition to the classifications in 2D groups in TABLE 1, we can also expand it to 3D systems through the corresponding mappings from layer groups to 3D space groups Litvin and Wike (2012). Each layer group is isomorphic to the factor group $G/T(1)$, where $G$ represent a 3D space group, and $T(1)$ is a 1D translation subgroup of $G$. It allows each layer group to correspond to a specific 3D space group, which represents the structure obtained when the layer group is extended along the $z$ direction. The corresponding space groups and required Wyckoff positions of the magnetic atoms for 3D materials are summarized in TABLE 3. For 3D materials with sufficiently large spacing between their mirror planes, one can also determine whether they possess decoupled spin and spatial symmetry using TABLE 3.
For 1D materials, their symmetry is typically described using rod groups, which possess translational symmetry along a single axis (marked as $z$-direction) with 3D spaceHahn et al. (1983). Based on the symmetry analysis above, we provide the TABLE 4, when magnetic materials belongs to the given rod groups and the magnetic atoms occupy the required Wyckoff positions, while their magnetic moments lie in the $xy$ plane, these materials could preserve spin-coplanar SSG symmetries described by Eq. (1) and Eq. (2). The TABLE 4 can also be expanded to identify the quasi-1D magnetic candidate materials. Since many magnetic materials exhibit quasi-1D characteristics, such as Sr3MIrO6(M = Li, Na, Zn, Cd) Segal et al. (1996), thereby the exploration of SSG symmetry in these materials is worthy of further investigation.
Acknowledgements
This work was financially supported by the National Natural Science Foundation of China (Grants No. 12188101, No. 12334007, and No. 12474233), the National Key R&D Program of China (Grant No. 2022YFA1403601), Innovation Program for Quantum Science and Technology (Grant No.2021ZD0301902, and No. 2024ZD0300101), Natural Science Foundation of Jiangsu Province (No. BK20233001, BK20243011).